the equation x(squared)+5x=0 has two solutions. whats the other?
Answer by priya(51)

THE given equation is x^2 + 5x = 0
x(x + 5) = 0(taking x as common )
we get 2 solutions,
that is x = 0 and x + 5 = 0
x = -5
the values are 0 and -5.
want to learn maths online..visit my site!!
priya